# Towards Imperceptible Watermarking Via Environment Illumination for Consumer Cameras

### v1

@@ 4.2 Human Sensitivity Evaluation
To assess the perceptibility of the proposed method, we evaluated human color perception under LED illumination optimized using the proposed spectra. For this evaluation, we used the lightbox setup shown in Fig. 4, which contained various colored objects for participants to observe.
@@
    Baseline Illumination Pattern: A conventional VLC approach that varies light intensity for data transmission. For comparison in this work, an intensity ratio of 1:0.97 with a white LED was selected to achieve the same pixel value difference as the optimized spectra.
@@
    Viewing Optimized Steady State Baseline: Participants first observed the objects under a steady, non-flickering optimized light for 10 s to understand the arrangement of objects.
@@
Illumination Type Count (Range) Rate
Static Optimized 0.133 (0-2) 2.67%
Optimized Flicker 0.467 (0-2) 9.33%
Static White 0.267 (0-2) 5.33%
Baseline Flicker 4.933 (4-5) 98.67%
[…]
This experiment involved 15 participants, with results summarized in Table 1. Flicker with an intensity ratio of 1:0.97 was expected to be imperceptible based on a simulated $\Delta_{\mathrm{E}}$ value of 0.608 using synthetic ColorChecker data. However, participants detected it with high accuracy (0.2839 for the proposed method). Under the baseline flicker condition, a near-universal detection rate of 98.67% was observed, confirming that this type of flicker was consistently noticeable to human observers.
In contrast, the optimized flicker condition achieved a low detection rate, with a mean of only 9.33%, suggesting that the optimized spectral modulation is largely imperceptible to humans. Meanwhile, static optimized and static white illuminations had even lower false-positive rates of 2.67% and 5.33%, likely due to minor fluctuations or momentary perceptions influenced by visual adaptation. These low rates confirm that static illuminations are indeed imperceptible, with occasional false positives attributable to natural perceptual variance.
The optimized flicker condition showed a slightly higher detection rate than false positives but remained effectively imperceptible to most participants. Since no participant detected the flicker in more than half of the five trials, it is reasonable to conclude that the optimized flicker is invisible to human observers.
Collectively, these results validate spectral optimization as a viable approach for low-frequency modulation in VLC systems, effectively circumventing human detection. Additionally, they confirm that conventional intensity modulation is unsuitable for low-frequency applications, as humans are highly sensitive to flicker in this range.
@@ 4.3 Evaluation of Decoding
Communication Accuracy in a Single Scene First, we evaluated the communication accuracy of the proposed method in a controlled environment using a single scene. We conducted long-duration recordings with three cameras (iPhone 13, Pixel 8 Pro, and Nikon D7200), each receiving 6400 bits of data. This setup established a baseline measure of communication accuracy for each camera and assessed whether the proposed method could achieve consistent performance across different devices. The results are shown in Fig. 7.
@@
Method Robustness in Diverse Environments Since practical applications involve diverse lighting conditions and scene compositions, additional tests were conducted to evaluate the proposed method’s robustness.
First, we used two smartphones to capture 128 bits of data in each of ten scenes to assess accuracy across different scenes. Examples of these scenes are shown in Fig. 8. This experiment evaluated the impact of scene diversity on communication accuracy and explored the range of environments where the proposed method could be effectively applied.
Additionally, to test robustness against variations in lighting conditions and camera angles, we adjusted the distance between the light source and the object to 88.9 cm, 76.2 cm, and 63.5 cm, positioning the smartphones at ten different angles for each distance. In each configuration, 128 bits of data were transmitted, allowing us to assess the impact of lighting and camera angles on communication accuracy.
@@
By employing dynamic thresholding, the proposed method demonstrated reliable signal recovery even in scenes with varying reflectance distributions and viewing angles. As shown in Fig. 8, the recovered signal scales differently across scenes due to reflectance variations, even after normalizing pixel values. However, dynamic thresholding ensures consistent signal recovery despite these scale differences, enabling accurate communication in diverse conditions.
@@ 5 Conclusion
Our findings suggest that spectral modulation from LED lighting holds significant potential for privacy protection and verification applications. By leveraging imperceptible spectral variations rather than intensity differences, our method enables nonintrusive, optically embedded watermarks that preserve the visual experience for human viewers while ensuring reliable detection by camera sensors.
Limitations and Future Directions
While our method considers the average pixel difference across frames, providing some robustness, it may still be sensitive to pixel fluctuations caused by object reflectance changes or additional ambient lighting. Future work will include a comprehensive evaluation of environmental impacts on detection accuracy and, if needed, algorithmic refinements to improve robustness against these variations.
Another limitation of our approach is its relatively low communication speed, capped at 15 fps, which may limit data transfer in time-sensitive applications. One potential solution is to increase the data rate by parallelizing communication across multiple wavelength channels, enabling independent data transmission on each channel.

### v2

@@ 4.2 Human Sensitivity Evaluation
Experimental Setup. To assess the perceptibility of the proposed method, we evaluated human color perception under LED illumination optimized using the proposed spectra. For this evaluation, we used the lightbox setup shown in Fig. 4, which contained various colored objects for participants to observe.
@@
    Baseline Intensity Pattern: A conventional VLC scheme that varies light intensity for data transmission (ratio 1:0.97).
To equate the sensor-level signal, the baseline ratio was tuned such that, under the same camera and exposure, the mean per-frame pixel-value difference between its two intensity levels matched that observed between the optimized spectra.
@@
    Viewing Steady-State Baseline: Participants first observed the objects under a steady optimized light for 10 sec to understand the arrangement of objects.
@@
Illumination Type Avg. detections (range) Detection rate
Optimized Steady 0.133 (0–2) 2.67%
Optimized Flicker 0.467 (0–2) 9.33%
Baseline Steady 0.267 (0–2) 5.33%
Baseline Flicker 4.933 (4–5) 98.67%
Table 1: Detection results. For each condition, 15 participants each performed five trials. “Avg. detections” is the average number of detections per participant over five trials (parentheses show the participant-wise min–max). “Detection rate (%)” equals (avg. detections / 5) $\times$ 100.
Results. We tested 15 participants (Table 1). Each participant was exposed to four illumination conditions: (i) Optimized Steady State; (ii) Optimized Flicker State ($\Delta E=0.2839$); (iii) Baseline Steady White Light; and (iv) Baseline Flicker White Light with an intensity ratio of 1:0.97 ($\Delta E=0.6080$). Although $\Delta E$ values in this range are typically regarded as below the threshold of noticeability for static color differences, in our low-frequency flicker experiment, the Baseline Flicker White Light condition (1:0.97) was detected by nearly all participants (98.67%), indicating that such small chromatic differences become clearly visible when presented as low-frequency temporal flicker.
By contrast, the optimized spectral modulation yielded a mean detection rate of only 9.33%; moreover, no participant reported it in more than half of the five trials, indicating that it was effectively imperceptible for most observers.
For the static controls, false-positive rates were lower still (2.67% for static optimized and 5.33% for static white), likely reflecting minor fluctuations or momentary adaptation. Taken together, these results support spectral optimization as a viable strategy for low-frequency modulation in imperceptible watermarking applications. At the same time, the findings underscore that conventional intensity modulation is unsuitable in this regime due to human flicker sensitivity.
@@ 4.3 Evaluation of Decoding
Communication Accuracy in a Single Scene. First, we evaluated the communication accuracy of the proposed method in a controlled environment using a single scene. We conducted long-duration recordings with three cameras (iPhone 13, Pixel 8 Pro, and Nikon D7200), each receiving 6400 bits, over several minutes per device. This setup established a baseline measure of communication accuracy for each camera and assessed whether the proposed method could achieve consistent performance across different devices. The results are shown in Fig. 7.
@@
Method Robustness in Diverse Environments. Since practical applications involve diverse lighting conditions and scene compositions, additional tests were conducted to evaluate the proposed method’s robustness.
First, we used two smartphones to capture 128 bits of data in each of ten scenes to assess accuracy across different scenes. Examples of these scenes are shown in Fig. 8. This experiment evaluated the impact of scene diversity on detection accuracy and explored the range of environments where the proposed method could be effectively applied.
Additionally, to test robustness against variations in lighting conditions and camera angles, we adjusted the distance between the light source and the object to 88.9, 76.2, and 63.5 cm (=35, 30, and 25 in), positioning the smartphones at ten viewing angles for each distance. In each configuration, 128 bits of data were transmitted, allowing us to assess the impact of lighting and camera angles on communication accuracy.
@@
Using dynamic thresholding, the proposed method demonstrated reliable signal recovery even in scenes with varying reflectance distributions and viewing angles. As shown in Fig. 8, the recovered signal scales differently across scenes due to reflectance variations, even after normalizing pixel values. Dynamic thresholding ensures consistent signal recovery despite these scale differences, enabling accurate communication in diverse conditions.
@@ 5 Conclusion
Our findings suggest that spectral modulation from LED lighting holds significant potential for privacy protection and verification applications. By leveraging imperceptible spectral variations rather than intensity differences, our method enables non-intrusive, optically embedded watermarks that preserve the visual experience for human viewers while ensuring reliable detection by camera sensors.
Limitations and Future Directions. While our analysis focuses on static objects, the same conclusion typically holds under small motions when the global image sum is stable, since $\sum(I_{t-1}-I_{t})=\sum I_{t+1}-\sum I_{t}$. As shown in the supplementary material, deviations arise when $\sum I_{t}$ changes—e.g., due to non-Lambertian/specular effects, ambient-light shifts, exposure/white-balance adjustments, or occlusions/parallax—which we leave for future evaluation and algorithmic refinements.
Another limitation of our approach is its relatively low communication speed, capped at 15 bps, which may limit data transfer in time-sensitive applications. One potential solution is to increase the data rate by parallelizing communication across multiple wavelength channels, enabling independent data transmission on each channel.